YoursTruly:

The table below shows the number of chapters in a book in relation to the number of typos found in the book. What is the equation for the line of best fit.

YoursTruly:

Screen%20Shot%202016-03-22%20at%203.08.33%20PM(1).png

YoursTruly:

A y=(4/5)x+(13/5) B y=8x+9 C y=3x+5

xxhero782xx:

i belive its B

xxhero782xx:

@theDeviliscoming